Comprehensive variance analysis review. Memlash Inc. manufactures 4 gigabyte flash drives that are compatible with a popular portable storage device. Memlash sells flash drives directly to computer retail chains and to direct marketing organizations that resell flash drives under their house brands. The flash drives retail for an average of S9.60 per unit, and compete with well-known brands that retail for between S12.00 and S14.40 per flash drive. Memflashs CFO has provided you with the following budgeted standards for the month of February 2015: Budgeted average wholesale selling price per unit Total direct material standard cost per drive Direct manufacturing labour Direct manufacturing labout standard cost per hour S18.00 Average labour productivity (drives per hout Direct marketing cost per unit Total fixed overhead The VP of marketing forecasts sales of 1,660,500 units for the month. On March 7, the VP of planning and control meets with the executive committee to discuss February results. He reports as follows: S 4.80 S 1.02 300 S 0.36 $1,080,000 Unit sales totalled 1,400,000 units. Actual average selling price declined to $4.86. Productivity dropped to 280 drives/hour; however, because of favourable market conditions, the actual direct materials cost per unit dropped to $0.94 Fixed costs came in $33,000 below All other costs were incurred at their standard rates. Required As the senior financial analyst, you are asked to calculate the following 1. Static-budget and actual operating income 2. Total static-budget variance. 3. Flexible-budget operating income. 4. Total flexible-budget variance 5. Total sales-volume variance 6. Rate and efficiency variances 7. What is the material-rate variance? What is the labour-rate variance? 8. What is the material-efficiency variance? What is the labout-efficiency variance? plan.
Answer #1

1. Static Budget for Feb 2015:

Sales quantity - 1660500 units

Revenue - $4.8 x 1660500 = $7,970,400

Direct Material Cost = $1.02 x 1660500 = $1693710

Direct Labor Cost = $18 x (1660500/300) = $99630

Direct Marketing Cost = $0.36 x 1660500 = $597,780

Total Budgeted Variable Costs = $2391120

Budgeted Fixed Costs = $1080,000

Budgeted Operating Income = $4,499,280

Actual Operating Income = ($4.86 - $0.94 - $18/280) x 1400,000 - $1047,000 = $4351000

2. Total Static Budget Variance = Actual - Budgeted Operating Income

= $4351000 - $4499280 = - $148,280

3. Flexible Budget Operating Income

Budgeted Revenue = $4.8 x 1400,000 = $6,720,000

Budgeted Direct Material Cost = $1.02 x 1400000 = $1428,000

Budgeted Direct Labor Cost = $18 x (1400000/300) = $84,000

Budgeted Fixed Costs = $1080,000

Flexible Budget Operating Income = $4128,000

4. Total flexible Budget Variance = Actual - Flexible budget Operating Income

=$4351000 - $4128000 = $223000

5. Total Sales Volume Variance = Actual - Budgeted Sales Volume

= 1400000-1660500 = - 260500

6. Rate and Efficiency Variances

Sales price variance = $4.86 - $4.8 = $0.06 (Positive)

Material Cost Variance = $0.94 - $1.02 = - $0.08 (Positive)

Efficiency Variance - Here only the labor efficiency variance is considered as efficiency of other processes like machine productivity, raw material consumption is assumed to be constant.

Labor efficiency variance = 280 - 300 = - 20 units / hr (Adverse)

7.Material Rate Variance = - $0.08, Labor rate Variance = 0

8. Material Efficiency variance = 0. Since it is mentioned that "All other costs were incurred at their standard rates" we have to assume that consumption of raw material per kg or finished goods remains the same as budgeted. Therefore material efficiency variance is zero.

Labor Efficiency variance = -20 / hr (Adverse). The budgeted efficiency was 300 units per hour but the actual efficiency was 280 units per hour and hence an adverse variance.